OpenStudy (anonymous):

Zoe factored the expression 16x - 28y + 24xy. Identify the term that is incorrect in her answer. 4(4x - 7y + 12y) A. 4 B. 4x C. 7y D. 12y

OpenStudy (zale101):

Combine like term inside the parenthesis

OpenStudy (anonymous):

How exactly?

OpenStudy (zale101):

First, work inside the parenthesis. Recall PEMDAS

OpenStudy (anonymous):

4 - 7 + 12 = Negative something? What am I missing.

OpenStudy (zale101):

\(4(4x \color{red}{-7y + 12y})\) -7y +12y are like terms, so add those up first.

OpenStudy (anonymous):

didn;'t know 7 was negative mb.

OpenStudy (zale101):

4x is a different term. You don't add those with y.

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

-7 + 12 = 5

OpenStudy (anonymous):

so 4x and 5y?

OpenStudy (zale101):

Yes. \(4(4x \color{red}{-7y + 12y})\) \(4(4x \color{red}{+5y})\)

OpenStudy (zale101):

Now, distribute the 4 in.

OpenStudy (anonymous):

LOL rhonda pls

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

which 4? the 4x or the 4 outside the parenthesis?

OpenStudy (zale101):

Ronda, that is incorrect and please don't give direct answers (even though is not the answer).

OpenStudy (anonymous):

Is it wrong to laugh at her..? XD

OpenStudy (zale101):

No. Anyways. Did you distribute the 4 into the parenthesis?

OpenStudy (anonymous):

I don't know how? would it be like 4 + 4x?

OpenStudy (anonymous):

I'm no human calculator XD

OpenStudy (zale101):

Are you familiar with distributing? Recall the distributive property: \(\Large a(b+c)=ab+ac\)

OpenStudy (anonymous):

Not exactly familiar with it. Would I do somethign like? 4+4x = 8x?

RhondaSommer (rhondasommer):

16x + -28y + 24xy Reorder the terms: 16x + 24xy + -28y Factor out the Greatest Common Factor (GCF), '4'. 4(4x + 6xy + -7y) Final result: 4(4x + 6xy + -7y) therefore; 12y would be the incorrect term

OpenStudy (anonymous):

Or would I do something like 4+4x = 8x and 4 + 5y = 9y?

OpenStudy (zale101):

Imagine a is 4, b is 4x, and c is 5y so \(4(4x+5y)=?\)

OpenStudy (anonymous):

I think I am getting it a little bit better.

OpenStudy (anonymous):

4b + 4c?

OpenStudy (zale101):

Actually, i read your question wrong :|

OpenStudy (anonymous):

o

OpenStudy (zale101):

Zoe factored the expression \(16x - 28y + 24xy. \) I thought we were just simplify this expression \(4(4x - 7y + 12y)\)

OpenStudy (zale101):

Can you factor this expression \(16x - 28y + 24xy\) ?

OpenStudy (zale101):

My apologies @AnimeWover

RhondaSommer (rhondasommer):

I factored it!

RhondaSommer (rhondasommer):

I told you what was wrong!

RhondaSommer (rhondasommer):

and i explained it!

OpenStudy (anonymous):

was rhona actually right?

OpenStudy (zale101):

@AnimeWover IS there a common factor in this expression 16x - 28y + 24xy? I doesn't mattr who's write, this site is not for direct answer but rather letting the asker do all the work while tutors make corrections.

RhondaSommer (rhondasommer):

16x + -28y + 24xy Reorder the terms: 16x + 24xy + -28y Factor out the Greatest Common Factor (GCF), '4'. 4(4x + 6xy + -7y) Final result: 4(4x + 6xy + -7y) 24y is incorrect

OpenStudy (anonymous):

Rhona strikes again! and common factor? What do u mean by that

RhondaSommer (rhondasommer):

it means that it can all be divided by one specific number

OpenStudy (anonymous):

Then yes. 4?

RhondaSommer (rhondasommer):

yep

RhondaSommer (rhondasommer):

and i meant 12y not 24y lol

OpenStudy (zale101):

@AnimeWover can you please show your work on why you think the answer is C? So if i said Rhonda is right, then that means you understood every concept in this question?

OpenStudy (anonymous):

What I would do is find a common factor to divide the numbers stated with?

OpenStudy (anonymous):

That is stated*

RhondaSommer (rhondasommer):

yes exactly!

RhondaSommer (rhondasommer):

most times, it is obvious and if there isnt one, then you can always just leave it simplified as it is

OpenStudy (anonymous):

Still Rhonda, Zale helped me by carefully explaining everything. @RhondaSommer You could learn a thing or two from @Zale101

OpenStudy (zale101):

Find the comon factor, compare it with \(4(4x - 7y + 12y)\) and see where is the mistake.

OpenStudy (zale101):

common*

OpenStudy (anonymous):

XD

OpenStudy (zale101):

makes sense? :)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

12y is incorrect

OpenStudy (anonymous):

it would be 6y

OpenStudy (anonymous):

24 / 4 = 6 12y is the incorrect term

RhondaSommer (rhondasommer):

6xy

OpenStudy (anonymous):

..?

RhondaSommer (rhondasommer):

look at the question. it is (xy) not just y thats one of the reasons it is wrong

RhondaSommer (rhondasommer):

brb...zale can you explain it to her? family stuff...

OpenStudy (zale101):

\(16x - 28y + 24xy\) 4 is the common factor \(4(4x - 7y + 6xy)\) compare this with \(4(4x - 7y + 12y) \)

OpenStudy (zale101):

Alright, take care @RhondaSommer !

OpenStudy (anonymous):

ohhh

OpenStudy (zale101):

everything seems similar except what?

OpenStudy (anonymous):

I kinda get it? I think

OpenStudy (zale101):

Is \(4(4x - 7y + 6xy)\) similar to \(4(4x - 7y + 12y)\)?

OpenStudy (zale101):

Yes or no?

OpenStudy (anonymous):

No. 6 has xy while the 12 has y

OpenStudy (zale101):

the first one has 6xy, the second one has 12y at the end. So 12y should not be there.

OpenStudy (anonymous):

true